LadyRavyRay:

The graph shows the number of servings in different amounts of frozen yogurt listed on a carton. Enter an equation that gives the number of servings y in x pints. Use only decimal form, if necessary. The equation is y =

LadyRavyRay:

jhonyy9:

look please on this graph y = 3 when x = 2 y = 6 when x = 4 y = 9 when x = 6 so what result from this ? 3 = 1,5 *2 6 = 1,5 *4 9 = 1,5 *6 hope helped understandably ? @XioGonz

XioGonz:

Yes Correct, I agree with jhony!

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Yes Correct, I agree with jhony! \(\color{#0cbb34}{\text{End of Quote}}\) ty

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Yes Correct, I agree with jhony! \(\color{#0cbb34}{\text{End of Quote}}\) ty \(\color{#0cbb34}{\text{End of Quote}}\) Np

draebrewer:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Yes Correct, I agree with jhony! \(\color{#0cbb34}{\text{End of Quote}}\) ty \(\color{#0cbb34}{\text{End of Quote}}\) Np \(\color{#0cbb34}{\text{End of Quote}}\) great

LadyRavyRay:

So...y = 9?